# Learning from Mistakes: Enhancing Harmful Meme Detection via Misjudgment Risk Patterns

### v1

@@ 4.2. Baselines
Specialized Approaches for Harmful Meme Detection. This category includes methods specifically designed or fine-tuned for this task. We consider two sub-types. The first are supervised methods that require task-specific training, including PromptHarm (Cao et al., 2022), Pro-Cap (Cao et al., 2023), and CapAlign (Ji et al., 2024). The second, which we refer to as MemeAgent (Lin et al., 2025), is a SOTA multi-agent framework for meme analysis that operates without requiring end-to-end training. For a fair comparison, we utilize the implementations of all these methods as provided in the MemeAgent study.
General-Purpose MLLMs. To demonstrate the effectiveness and generalizability of our approach, we establish baselines using a range of SOTA, general-purpose MLLMs. This includes Qwen2.5-VL-32B-Instruct (Bai et al., 2025), InternVL3.5-20B-A4B (Wang et al., 2025), InternVL3-14B (Zhu et al., 2025), LLaVA-1.6-13B [^3] The baseline performance for LLaVA-1.6-13B is sourced from the MemeAgent (Lin et al., 2025) study for consistency.  (Liu et al., 2024), GPT-4o (Hurst et al., 2024), and GPT-4o-mini (Hurst et al., 2024). We evaluate their native performance in a zero-shot CoT setting.
@@ 5. Results and Analysis
Table 3. The results of ablation studies of the PatMD (%).
@@
PatMD 76.95 75.06 77.80 77.60 63.26 57.18 80.03 79.88 66.70 64.99 72.95 70.94
w/o tree_matching 74.26 71.36 77.30 77.19 62.25 55.37 78.26 78.18 64.11 62.22 71.24(-1.71) 68.86(-2.08)
w/o img_txt_rerank 76.19 74.29 76.98 76.73 62.05 56.32 78.95 78.78 66.13 64.49 72.06(-0.89) 70.12(-0.82)
w/o txt_retrieval 75.70 73.74 77.34 77.29 62.48 56.39 78.04 77.89 64.35 62.26 71.58(-1.37) 69.51(-1.43)
w/o img_retrieval 75.64 73.67 77.10 76.83 62.99 55.77 79.35 79.20 65.10 62.99 72.04(-0.91) 69.69(-1.25)
random retrieval 73.07 69.83 74.54 74.54 61.89 54.15 75.06 74.80 63.55 61.00 69.62(-3.33) 66.86(-4.08)
@@ 5.1. Performance of PatMD (RQ1)
To evaluate the effectiveness of our proposed approach, we compare the performance of PatMD with the baselines and general-purpose MLLMs (i.e., Vanilla-MLLM in the table), as shown in Table 2.
Overall Performance. The results demonstrate that PatMD consistently improves the performance of specialized baselines and general-purpose MLLMs, with an average of 8.30% improvement in F1-scores and an average of 7.71% improvement in accuracy. First, when compared against specialized harmful meme detection baselines, our method shows strong performance. For instance, when applied to GPT-4o and the open-source InternVL3-14B, our framework achieves overall F1-scores of 74.67% and 70.94% respectively, both outperforming the strongest baseline, MemeAgent (65.19%). Second, the framework exhibits strong generalization across different MLLMs. It improves the overall F1-score of GPT-4o by 5.24% and that of GPT-4o-mini by 7.39%. The impact on open-source models is also notable. InternVL3-14B experiences an 11.50% F1-score increase, surpassing the vanilla GPT-4o (70.94% vs. 69.43%). Similarly, LLaVA-1.6-13B, a model of comparable scale, receives an F1-score boost of 7.85%. Furthermore, the framework proves effective even for specialized architectures such as InternVL3.5-20B-A4B, a Mixture-of-Experts (MoE) model activating only 4B parameters during inference, achieving an overall F1 improvement of 3.49%. These results validate that PatMD is an effective and model-agnostic approach capable of enhancing the detection performance of MLLMs across different scales and architectures.
Task-Specific Analysis. A detailed analysis of task-specific results further reveals the breadth and depth of the improvements. The framework’s impact is particularly evident in Sarcasm and Misogyny detection. Notably, for Sarcasm, PatMD yields substantial improvements across models. For instance, InternVL3-14B’s F1-score improves by 29.54%, and GPT-4o-mini’s F1-score increases by 30.81%. This highlights PatMD’s ability to help models interpret the incongruity between text and image—a common characteristic of sarcastic memes. In Misogyny detection, similar improvements are observed. Qwen2.5VL-32B’s F1-score, for example, increases by 17.95%. This indicates that by retrieving historical samples with similar stereotypical patterns, our framework effectively equips models to identify targeted, gender-based harm. Beyond these tasks, PatMD also demonstrates effectiveness across other domains. In Offensiveness detection, most models show positive gains, such as the 6.67% F1-score increase for GPT-4o-mini. For Harmfulness, a general positive trend is also observed, with Qwen2.5VL-32B gaining 9.89% in F1-score. These widespread gains underscore the framework’s ability to provide task-relevant reasoning support.
@@ 5.2. Ablation Study (RQ2)
To evaluate the effectiveness of the key components in our PatMD, we conduct a series of ablation studies. We evaluate the impact on overall performance by removing or replacing key components of our approach, including: the structural matching of hierarchical harm tree in Section 3.2.2 (w/o tree_matching), the image and text reranking in Section 3.2.2 (w/o img_txt_rerank), the text/image retrieval in Section 3.2.1 (w/o txt/img_retrieval), and replacing the bipartite candidate retrieval of Section 3.2.1 with random retrieval (random retrieval). The experimental results are presented in Table 3.
As shown in Table 3, the complete PatMD approach achieves the best performance in both accuracy (Acc) and F1 score, demonstrating the integrity and synergistic effect of our design. Among all variants, removing the structural matching of hierarchical harm tree (w/o tree_matching) leads to one of the most significant performance drops, with the F1 score decreasing by 2.08% (from 70.94% to 68.86%). This strongly demonstrates that our proposed hierarchical harm tree is crucial for the model to understand and judge the deep-seated risks in complex memes. Removing the image and text reranker (w/o img_txt_rerank), which is used to refine content similarity, also results in a performance decline. This indicates that a more precise assessment of content relevance is essential besides performing the structure matching. Furthermore, performance drops when the text retrieval (w/o txt_retrieval) and image retrieval (w/o img_retrieval) components are removed individually, validating the necessity of leveraging both modalities for retrieval, as relying on a single modality can miss key similar samples. Finally, compared to random retrieval (random retrieval), our method shows a substantial performance advantage, with a F1 score more than 4% higher. This highlights that accurately retrieving highly relevant historical memes from the knowledge base is vital to ensuring the overall effectiveness of the framework.
@@ 6. Discussion
Figure 4. The case study of PatMD on four examples.
@@ 6.2. Human Evaluation
We instruct three human evaluators, each with over five years of experience in digital content moderation, to rate these patterns on a 5-point Likert scale (from 1: poor to 5: excellent) based on four key criteria: relevance, correctness, actionability, and uniqueness. The evaluation results, as illustrated in Figure 5, clearly indicate that the majority of ratings are concentrated between 4 and 5. This suggests that the misjudgment risk patterns generated by GPT-4o are generally well-perceived and align closely with human judgment on potential risks.
@@ 7. Conclusion
In this paper, we propose PatMD, an approach that enhances harmful meme detection by learning from potential misjudgment risks. It first constructs a knowledge base, where each meme is deconstructed into a misjudgment risk pattern that captures why and how the MLLM may misinterpret its content. For a given target meme, PatMD retrieves the semantically similar memes, and the relevant patterns are used to dynamically guide the MLLM’s reasoning, proactively guiding the MLLMs to avoid known misjudgment pitfalls. Experiments show that PatMD significantly outperforms the baselines, achieving an average of 7.26% F1-score improvement.
@@ 18  $\text{TopK\_Cases}\leftarrow$ Select top $K$ items from ScoredList;
20 $P_{Hateful}\leftarrow\textsc{ExtractPattern}(\text{TopK\_Cases},\text{'hateful'})$;
21  $P_{Benign}\leftarrow\textsc{ExtractPattern}(\text{TopK\_Cases},\text{'benign'})$;

### v3

@@ 4.2. Baselines
Specialized Approaches for Harmful Meme Detection.
This category includes methods specifically designed or fine-tuned for this task. We consider two sub-types. The first are supervised methods that require task-specific training, including PromptHarm (Cao et al., 2022), Pro-Cap (Cao et al., 2023), and CapAlign (Ji et al., 2024). The second subtype comprises training-free frameworks, represented by MemeAgent (Lin et al., 2025), a SOTA multi-agent framework for meme analysis. We strictly adhere to the rigorous evaluation protocol and implementations established by the MemeAgent study to ensure a fair comparison.
General-Purpose MLLMs. To demonstrate the effectiveness and generalizability of our approach, we establish baselines using a range of SOTA, general-purpose MLLMs. This includes Qwen2.5-VL-32B-Instruct (Bai et al., 2025), InternVL3.5-20B-A4B (Wang et al., 2025), InternVL3-14B (Zhu et al., 2025), LLaVA-1.6-13B [^2] The baseline performance for LLaVA-1.6-13B is sourced from the MemeAgent (Lin et al., 2025) study for consistency.  (Liu et al., 2024), GPT-4o (Hurst et al., 2024), and GPT-4o-mini (Hurst et al., 2024).
@@ 5. Results and Analysis
Table 3. The results of ablation studies between the PatMD and RAG-oriented variants (%).
@@
w/o Structural Reranking 74.26 71.36 77.30 77.19 62.25 55.37 78.26 78.18 64.11 62.22 71.24(-1.71) 68.86(-2.08)
w/o Semantic Reranking 76.19 74.29 76.98 76.73 62.05 56.32 78.95 78.78 66.13 64.49 72.06(-0.89) 70.12(-0.82)
Image-only RAG 75.70 73.74 77.34 77.29 62.48 56.39 78.04 77.89 64.35 62.26 71.58(-1.37) 69.51(-1.43)
Text-only RAG 75.64 73.67 77.10 76.83 62.99 55.77 79.35 79.20 65.10 62.99 72.04(-0.91) 69.69(-1.25)
Random RAG 73.07 69.83 74.54 74.54 61.89 54.15 75.06 74.80 63.55 61.00 69.62(-3.33) 66.86(-4.08)
@@ 5.1. Performance of PatMD (RQ1)
To evaluate the effectiveness of our proposed approach, we compare the performance of PatMD with the baselines and general-purpose MLLMs (i.e., MLLM in the table), as shown in Table 2.
Overall Performance. The results demonstrate that PatMD consistently improves the performance of specialized baselines and general-purpose MLLMs, with an average of 8.30% improvement in F1-scores and an average of 7.71% improvement in accuracy. First, when compared against specialized harmful meme detection baselines, our method shows strong performance. For instance, when applied to GPT-4o and the open-source InternVL3-14B, our framework achieves overall F1-scores of 74.67% and 70.94%, respectively, both outperforming the strongest baseline, MemeAgent (65.19%). Second, the framework exhibits strong generalization across different MLLMs. It improves the overall F1-score of GPT-4o by 5.24% and that of GPT-4o-mini by 7.39%. The impact on open-source models is also notable. InternVL3-14B experiences an 11.50% F1-score increase, surpassing the GPT-4o (70.94% vs. 69.43%). Similarly, LLaVA-1.6-13B, a model of comparable scale, receives an F1-score boost of 7.85%. Furthermore, the framework proves effective even for specialized architectures such as InternVL3.5-20B-A4B, a Mixture-of-Experts (MoE) model activating only 4B parameters during inference, achieving an overall F1 improvement of 3.49%. These results validate that PatMD is an effective and model-agnostic approach capable of enhancing the detection performance of MLLMs across different scales and architectures.
Task-Specific Analysis. A detailed analysis of task-specific results further reveals the breadth and depth of the improvements. The framework’s impact is particularly evident in Sarcasm and Misogyny detection. Notably, for Sarcasm, PatMD yields substantial improvements across models. For instance, InternVL3-14B’s F1-score improves by 29.54%, and GPT-4o-mini’s F1-score increases by 30.81%. This highlights PatMD’s ability to help models interpret the incongruity between text and image—a common characteristic of sarcastic memes. In Misogyny detection, similar improvements are observed. Qwen2.5VL-32B’s F1-score, for example, increases by 17.95%. This indicates that by retrieving historical samples with similar stereotypical patterns, our framework effectively equips models to identify targeted, gender-based harm.
Beyond these tasks, PatMD also demonstrates effectiveness across other domains. In Offensiveness detection, most models show positive gains, such as the 6.67% F1-score increase for GPT-4o-mini. For Harmfulness, a general positive trend is also observed, with Qwen2.5VL-32B gaining 9.89% in F1-score. These widespread gains underscore the framework’s ability to provide task-relevant reasoning support.
@@ 5.2. Ablation Study (RQ2)
To evaluate the effectiveness of the key components in our PatMD, which covers all the RAG’s framework in Figure 2. We conduct a series of ablation studies, which evaluate the impact on PatMD’s performance by removing or replacing key components (related to RAG paradigm) of our approach, including: the structural matching of hierarchical harm tree in Section 3.2.2 (w/o Structural Reranking), the image and text reranking in Section 3.2.2 (w/o Semantic Reranking), the text/image retrieval in Section 3.2.1 (w/o Image/Text-only RAG), and replacing the bipartite candidate retrieval of Section 3.2.1 with random retrieval (Random RAG).
As shown in Table 3, the complete PatMD approach achieves the best performance in both accuracy and F1 score, demonstrating the integrity and synergistic effect of our design. Among all variants, removing the structural matching of the hierarchical harm tree (w/o Structural Reranking) leads to one of the most significant performance drops, with the F1 score decreasing by 2.08% (from 70.94% to 68.86%). This strongly demonstrates that our proposed hierarchical harm tree is crucial for the model to understand and judge the deep-seated risks in complex memes. Removing the image and text reranker (w/o Semantic Reranking), which is used to refine content similarity, also results in a performance decline. This indicates that a more precise assessment of content relevance is essential, besides performing the structure matching. Furthermore, performance drops when the text retrieval (w/o Image-only RAG) and image retrieval (w/o Text-only RAG) components are removed individually, validating the necessity of leveraging both modalities for retrieval, as relying on a single modality can miss key similar samples. Finally, compared to random retrieval (Random RAG), our method shows a substantial performance advantage, with a F1 score more than 4% higher. This highlights that accurately retrieving highly relevant historical memes from the knowledge base is vital to ensuring the overall effectiveness of the framework.
@@
5.3. Generality and Robustness of PatMD (RQ3)
Table 4. Performance of PatMD across two evaluation settings (%): generalization on unseen memes, and robustness evaluation (original vs. adversarial) on the Hatefulness task. Values in parentheses indicate the performance drop compared to the original memes.
Approach Unseen Memes Robustness Evaluation
Original Memes Adversarial Memes
Acc F1 Acc F1 Acc F1
InternVL3-14B 67.00 66.51 68.25 68.24 67.75 (-0.50) 67.70 (-0.54)
+PatMD 72.60 71.78 76.95 75.06 75.54 (-1.41) 73.34 (-1.72)
$\Delta$ (Improvement) +5.40+5.40 +5.27+5.27 +8.70+8.70 +6.82+6.82 +7.79+7.79 +5.64+5.64
To further evaluate the generality and robustness of our proposed framework, we conduct additional experiments on out-of-distribution (unseen) and adversarial memes. We select InternVL3-14B as the backbone model for these evaluations.
Generality: Performance on Unseen Memes. To evaluate whether PatMD can generalize to entirely new topics without updating its knowledge base, we evaluate it on a newly curated set of 500 recent political memes, crawled from social media platform X spanning from April 1, 2025, to July 1, 2025 (detailed in Appendix), which are strictly excluded from the previous training and testing datasets. As shown in Table 4, relying solely on the existing misjudgment risk patterns retrieved from the original knowledge base, PatMD achieves an accuracy of 72.60% and an F1-score of 71.78%. This represents a significant improvement of 5.40% in accuracy and 5.27% in F1-score compared to the InternVL3-14B model. This substantial gain demonstrates that the misjudgment risk patterns extracted by PatMD capture the underlying cognitive logic of harm manifestation (e.g., rhetorical devices and implicit analogies) rather than merely memorizing dataset-specific content. Consequently, these misjudgment risk patterns can be effectively migrated to unseen domains and guide the model in detecting these new memes.
Robustness: Performance on Adversarial Memes. To evaluate the robustness of PatMD against adversarial attacks, we apply the SaltPepper-I-High attack (Aggarwal et al., 2023; Mei et al., 2025), which injects random white and black pixels across the original memes in a manner that does not compromise the overall perception of semantic content.
Under this high-intensity visual perturbation, PatMD achieves an accuracy of 75.54% and an F1-score of 73.34%. While this reflects a slight sensitivity to extreme noise compared to original images, such behavior can be attributed to its underlying mechanism: our approach relies on extracting fine-grained visual details for structural harm deconstruction, which are naturally more susceptible to severe pixel-level noise than superficial features. Nevertheless, despite this minor degradation, PatMD still significantly outperforms the vanilla model under the same adversarial conditions, delivering an improvement of 7.79% in accuracy and 5.64% in F1-score. This demonstrates that the decision-centric guidance provided by PatMD remains highly robust and effective even in challenging adversarial environments.
@@ 6.1. Case Study
Figure 4. The case study of PatMD.
To intuitively illustrate how our framework guides the reasoning process of MLLMs, we present a representative case study in Figure 4. Further analyses of diverse scenarios, such as identifying implicit stereotypes and navigating subjective satire, are provided in the Appendix  D.”
Accurate Sarcasm Detection through Contextual Incongruity. Figure 4 illustrates a meme using an absurd panda analogy for an anti-racist message. To prevent MLLMs from interpreting the lighthearted imagery literally, our framework retrieves critical misjudgment risk patterns. These patterns warn that irony is easily missed when “text contradicts absurd imagery” and emphasize its dependence on “sociopolitical context.” Guided by these insights, the MLLM’s reasoning (Output Thought) successfully captures the incongruity between the serious racial issue and the visual metaphor, accurately classifying it as “Sarcastic.” This case exemplifies how our decision-centric approach proactively mitigates cognitive blind spots. By explicitly injecting risk patterns into the reasoning chain, the framework not only achieves accurate classification but also yields a highly transparent and interpretable decision logic.
@@ 6.2. Human Evaluation
We instruct three human evaluators, each with over five years of experience in digital content moderation, to rate these patterns on a 5-point Likert scale (from 1: poor to 5: excellent) based on four key criteria: relevance, correctness, actionability, and uniqueness. The distribution of average scores of these criteria, as illustrated in Figure 5, indicates that the majority of ratings are concentrated between 4 and 5. This suggests that the misjudgment risk patterns generated by GPT-4o are well-recognized and align closely with human judgments of potential risks.
6.3. Limitations
Inference Latency. PatMD introduces an additional inference latency of approximately 0.7s per meme compared to the vanilla MLLM. This overhead stems from the retrieval and reranking of risk patterns, as well as processing the augmented context. However, this slight delay is a worthwhile trade-off for the substantial 8.30% F1-score improvement, particularly in scenarios prioritizing detection accuracy over strict real-time processing.
Dependency on MLLM Reliability. The automated pattern elicitation relies on the underlying MLLM, which is susceptible to hallucinations. While we mitigated this in our current knowledge base through human evaluation and manual correction of erroneous patterns, the framework’s scalability remains bounded by the MLLM’s reliability. Future work could explore leveraging advanced reasoning models or multi-agent consensus frameworks for automated pattern verification, thereby reducing human intervention and enhancing the robustness of the knowledge base.
@@ 7. Conclusion
In this paper, we propose PatMD, a cognitive-guided approach that detects harmful memes by learning from and proactively mitigating potential misjudgment risks. It first constructs a knowledge base, where each meme is deconstructed into a misjudgment risk pattern that captures why and how the MLLM may misinterpret its content. For a given target meme, PatMD retrieves the semantically similar memes, and the relevant patterns are used to dynamically steer the MLLM’s reasoning, proactively guiding the MLLM to avoid known misjudgment pitfalls. Experiments show that PatMD significantly outperforms the baselines, achieving an average of 8.30% F1-score improvement, while demonstrating generality and robustness on unseen and adversarial memes.
@@ 18  $\text{TopK\_Cases}\leftarrow$ Select top $K$ items from ScoredList;
20 $P_{Hateful}\leftarrow\textsc{ExtractPattern}(\text{TopK\_Cases},\text{`hateful'})$;
21  $P_{Benign}\leftarrow\textsc{ExtractPattern}(\text{TopK\_Cases},\text{`benign'})$;
@@ Appendix C Prompt for Harmful Meme Detection
Appendix D Case Study
Figure 7. The case study of PatMD on four examples.
Appendix E Details of the Unseen Meme Dataset
To rigorously evaluate the out-of-distribution (OOD) generalization capability of PatMD(as discussed in RQ3), we curated a new dataset comprising 500 recent political memes. This section details the data collection, annotation process, and statistical distribution of this unseen dataset.
Data Collection and Thematic Scope.
The dataset was sourced from the social media platform X (formerly Twitter) spanning from April 1, 2025, to July 1, 2025. This specific timeframe ensures that the collected memes encapsulate recent political events and emerging rhetorical trends that are strictly absent from our primary training and testing sets. To comprehensively evaluate the model’s robustness across diverse contexts, the curated memes cover a wide array of highly debated topics, including: (1) U.S. partisan politics and macro-ideologies; (2) foreign policy and international conflicts; (3) gender issues, LGBTQ+ rights, and family values; (4) gun rights and gun control; and (5) government credibility, economic policies, and social welfare.
Annotation Process.
The entire annotation process was strictly guided by a unified definition of harmful content (Lin et al., 2025): “speech or material that mocks or ridicules a targeted person or organization, or has the potential to cause emotional discomfort to any individual, politician, celebrity, or the general public.” Based on this explicit criterion, we adopted a robust model-in-the-loop annotation pipeline to balance efficiency and accuracy. Initially, an ensemble of multiple SOTA MLLMs was prompted with this definition to generate preliminary labels for each meme, followed by a majority voting mechanism to determine the consensus prediction. Subsequently, a rigorous manual review phase was conducted by human experts. These annotators verified the labels generated by the models against the predefined criteria, corrected any misclassifications (particularly for memes involving implicit satire or complex cultural contexts), and finalized the categorization.
Dataset Statistics.
The final dataset consists of 500 annotated memes, comprising 300 (60%) “Benign” samples and 200 (40%) “Harmful” samples (see Figure 8 for examples). We utilize this dataset to evaluate whether PatMD can effectively transfer its learned misjudgment risk patterns to novel, unseen political domains, rather than merely relying on prior knowledge specific to the dataset.
Figure 8. Examples of unseen memes.